### v1

@@ 5 Experiments
#authors #papers $\operatorname{obj}({\mathcal{S}}_{\text{greedy}})$ $\operatorname{obj}^{\prime}({\mathcal{S}}_{\text{greedy}})$ $\operatorname{obj}({\mathcal{S}}_{\text{random}})$ $\operatorname{obj}^{\prime}({\mathcal{S}}_{\text{random}})$
ICLR 2021 8875 2964 5242 1365 2515 1132
ICLR 2022 10583 3328 6123 1574 2966 1324
ICLR 2023 15372 4881 9764 2489 4600 2045
@@
In Table 4, we summarize the basic statistics of ICLR 2021-2023 that we collected as well as the partitions determined from their ownership sets. There is a clear growing trend in conference size in terms of submission and author number. Moreover, from both partition objective functions, the greedy partition is a clearly better choice than the randomly generated partition, though we are unable to efficiently find the exact optimal partition and empirically examine the greedy approximation ratio.
(a) ICLR 2021
(b) ICLR 2022
(c) ICLR 2023
Figure 3: The loss (normalized as the percentage change, $\frac{\text{model}-\text{baseline}}{\text{baseline}}$) of Isotonic Mechanism under varying level of review noise $\sigma$ on ICLR.
(a) ICLR 2021 Oral
(b) ICLR 2022 Oral
(c) ICLR 2023 Oral
(d) ICLR 2021 Spotlight
(e) ICLR 2022 Spotlight
(f) ICLR 2023 Spotlight
(g) ICLR 2021 Poster
(h) ICLR 2022 Poster
(i) ICLR 2023 Poster
Figure 4: The percentile precision of Isotonic Mechanism under varying level review noise $\sigma$ on ICLR.
@@
We compare the empirical performance of the Isotonic Mechanism with different baselines at different review noise levels. In Figures 3 and 4, we can see that the proposed Isotonic Mechanism can mitigate a substantial amount of noise from the review score and improve the precision metric over the baseline approach. The percentage loss change in Figure 3 suggests our proposed method is able to mitigate about $10-20\%$ amount of review noise in the past three years of ICLR. In addition, the performance of the Isotonic Mechanism steadily improves as the conference size grows — an encouraging sign for the applicability of the Isotonic Mechanism. Notably, the improvement on precision@30% suggests that $5\%\sim 15\%$ of the mistakenly rejected papers can be rectified to acceptance through the proposed score calibration method. Meanwhile, it is also intriguing to understand the degrading performance on precision@1.5%. This pattern is consistent in the past three years of ICLR. The potential reason may be twofold. On the one hand, the best papers may require more community critics rather than the authors’ self-evaluation, making the Isotonic Mechanism less applicable. On the other hand, it may attribute to the limitation of our data generation process, as we simulate the paper quality purely based on its review score, which is not necessarily a good indicator of a top paper being selected as “oral”.
Figure 5: The loss of Isotonic Mechanism under different levels of owners’ perception noise with varying number of owners per partition block and different review noise $\sigma=$ 1 (left), 2 (middle), 4 (right) on a $7$-level binary tree (top) and $11$-level ternary tree (bottom).
@@
Moreover, we seek to understand the structural influence of partition on the informativeness and robustness of the Isotonic Mechanism. An immediate observation is that the Isotonic Mechanism under the greedy partition outperforms that under the random partition, which matches our hypothesis in the theoretical analysis in Section 4. We also examine the empirical tradeoff of robustness and informativeness through a simulation of hierarchical ownership, given that the authors have a certain level of perception noises on their item qualities. We construct a $11$-level binary tree with $2^{11}$ items and $12^{11}-$ 1 owners, as well as a $7$-level ternary tree with $3^{7}$ items and $(3^{7}-1)/2$ owners. For the intuition, the partition at one extreme is to have only a single block of all items owned by the root node (owner), and there is only one owner we can elicit from. At the other extreme, the partition is to have one block for the set of items owned by each leaf node, and we can elicit from the owners at this leaf node and all its ancestor nodes. More generally, there exist $L$-strong partitions with $L$ from 1 to 11 in the binary tree, or from 1 to 7 in the ternary tree. In Figure 5, we can observe a clear trend of tradeoff: in the case of low noise (blue curve), the large block size is preferred to have more pairwise information, while in the case of high noise (red curve), the small block size is preferred to have more owners per block (larger $L$). In cases of medium noise (green or orange curve), the optimal block size is a careful balance of the block size and $L$-strongness. The trend showcases the importance of eliciting from multiple owners, e.g., to enforce the $L$-strong partition constraint, though the parameters shall be tailored carefully for the noise level in real-world applications.
6 Proofs
In this section, we prove Theorems 1 and 2 as well as some technical lemmas.
6.1 Proof of Theorem 1
@@
holds for all convex functions $h$ if and only if $\bm{a}\succeq\bm{b}$.
Proof of Theorem 1.
We first show that truthful reporting forms a Bayes-Nash equilibrium. Pick any owner and let its utility function be $\mathrm{U}(\widehat{\bm{R}})=\sum_{i=1}^{n}{U}(\widehat{R}_{i})$. Suppose that the other $m-1$ owners are already truth-telling and without loss of generality we can index them from $1$ to $m-1$. With slight abuse of notation, we let $\widehat{\bm{R}}^{j}(\pi)$ be the random variable for the item quality estimates using owner $j$’s reported ranking $\pi$ w.r.t. The randomness of review noise $\bm{z}$, i.e, $\mathop{\mathbf{Pr}}[\widehat{\bm{R}}^{j}(\pi)=\widehat{\mathrm{R}}^{j}(\pi;\bm{R}+\bm{z})]=\mathop{\mathbf{Pr}}[\bm{z}]$. Suppose that the first $m-1$ estimates use the true ranking and the true ranking $\pi^{\star}$ is the identity permutation for simplicity. It is clear that the item quality estimates with all $m-1$ owners’ truthfully reported rankings are all identical,
@@
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\alpha^{m}\widehat{\bm{R}}^{m}(\pi)+(1-\alpha^{m})\widehat{\bm{R}}(\pi^{\star}))\right]=\mathop{\mathbf{E}}_{\bm{z}}\left[\sum_{\rho}\mathrm{U}\left(\alpha^{m}\widehat{\mathrm{R}}^{m}(\pi;\rho\circ\bm{z})+(1-\alpha^{m})\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})\right)\right].$
@@
Let $\bm{a}^{+}$ be the projection of $\bm{a}$ onto the standard isotonic cone $\{\bm{a}|a_{1}\geq a_{2}\geq\dots\geq a_{n}\}$. Recall that we assume without loss of generality that $R_{1}\geq R_{2}\dots\geq R_{n}$. Following from the coupling argument in the proof of Theorem 2 in [23], we have $\mathrm{R}(\pi^{\star};\bm{z})=(\bm{R}+\bm{z})^{+}$ and $\mathrm{R}(\pi;\bm{z})=\pi^{-1}\circ(\pi\circ\bm{R}+\pi\circ\bm{z})^{+},\forall\bm{z}\in\mathbb{R}^{n}$. So we can rewrite the inequality (3) as,
@@
The first inequality is due to Lemmas 1 and 2, as $\pi^{-1}$ is some permutation on $\bm{b}^{+}=(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}$. The second inequality follows from Lemmas 1, 3 as well as Lemma 6.5 of [23], which shows that $(\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\succeq(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+},\forall\rho$. The third inequality follows from Lemma 1 and 2 where for each distinct permutation $\rho_{j}$, $(1-\alpha^{m})(\bm{R}+\rho_{j}\circ\bm{z})^{+}_{i},\alpha^{m}(\bm{R}+\rho_{j}\circ\bm{z})^{+}_{i}$ are the $j$-th entry of $\bm{a}^{+},\bm{b}^{+}\in\mathbb{R}^{n!}$ and $\pi$ is some permutation on $\bm{b}^{+}$. The equalities are by the definition of the utility function and its linearity.
@@
It now remains to prove that truthful NE gives every owner the highest equilibrium utility among all possible equilibria of the game. Again, we pick arbitrary owner $j$. Under truthful NE, the owner’s utility is $\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\widehat{\bm{R}}(\pi^{\star}))\right]$, where $\widehat{\bm{R}}(\pi^{\star})$ is the item quality estimate under the truthful ranking. We compare it to any (possibly mixed) Nash equilibrium (see Definition 1). Consider any pure strategy profile in this mixed Nash equilibrium, $\bm{\pi}=(\pi^{1},\pi^{2},\dots,\pi^{m})\sim\bm{x}=(x^{1},x^{2},\dots,x^{m})$. This sampled profile of reported ranking corresponds to the weighted average of quality estimates, $\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}(\pi^{j})$. We can see that the owner’s expected utility under this pure strategy profile is no larger than that under the truthful strategy profile,
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}\left(\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}(\pi^{j})\right)\right]$ $\leq\sum_{j=1}^{m}\alpha^{j}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\widehat{\bm{R}}(\pi^{j}))\right]$ (4)
$\leq\mathop{max}_{j\in[m]}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\widehat{\bm{R}}(\pi^{j}))\right]$
$\leq\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\widehat{\bm{R}}(\pi^{\star}))\right],$
@@
where the first inequality is due to Jensen’s inequality and linearity of expectation, the second inequality is by the property of arithmetic mean, last inequality is due to Theorem 6.1. Hence, the owner’s expected utility under this Nash equilibrium, taking an expectation over the distribution of pure strategy profile, $\mathop{\mathbf{E}}_{\bm{\pi}\sim\bm{x}}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}^{j})\right]\leq\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\widehat{\bm{R}}(\pi^{\star}))\right]$, is no larger than the expected utility under the truthful NE. ∎
Remark 2.3.

### v3

@@ 5.1 Experiment Setups: Datasets, Baselines and Metrics
#authors #papers $\operatorname{obj}({\mathcal{S}}_{\text{greedy}})$ $\operatorname{obj}({\mathcal{S}}_{\text{random}})$ $\operatorname{obj}^{\prime}({\mathcal{S}}_{\text{greedy}})$ $\operatorname{obj}^{\prime}({\mathcal{S}}_{\text{random}})$
ICLR 2021 8875 2964 5242 2515 1365 1132
ICLR 2022 10583 3328 6123 2966 1574 1324
ICLR 2023 15372 4881 9764 4600 2489 2045
@@
5.2 Experiment Results
Given the two rich sets of real world peer review data, we empirically study the following three questions about the performance of our proposed mechanism.
@@ 5.2 Experiment Results
Q1: How does the mechanism’s calibration quality change under different levels of noise? Here we use the ICLR dataset where we can control the noise level in our simulation of ground-truth score. Figure 5 compares the empirical performance of the Isotonic Mechanisms with different baselines at different review noise levels. We can see that the proposed Isotonic Mechanism can mitigate a substantial amount of noise from the review score and improve the precision metric over the baseline approach. The percentage MSE loss in Figure 5 suggests our proposed method is able to mitigate about $10-20\%$ amount of review noise in the past three years of ICLR. In addition, the performance of the Isotonic Mechanisms steadily improves as the conference size grows — an encouraging sign for the applicability of the Isotonic Mechanism. This performance is also reflected in determining accepted paper (i.e., top 30%) as illustrated in the lower panel of Figure 5. It suggests that $5\%\sim 15\%$ of the mistakenly rejected papers can be rectified to acceptance through the proposed score calibration method. [^13] In Appendix F, we plot how MSE improvement influences the accuracy of more stringent paper selections, e.g., spotlight (top 5%) and oral (top 1.5%) presentations. The performance has similar trend of improvement, hence is omitted here to avoid repetition.
(a) MSE on ICLR 2021
(b) MSE on ICLR 2022
(c) MSE on ICLR 2023
(d) Accept Accuracy
(e) Accept Accuracy
(f) Accept Accuracy
Figure 5: Plots of mechanism performances under varied noise level $\sigma$. Upper: The mean square error (MSE) loss of score estimation (normalized as the percentage change, $\frac{\text{model}-\text{baseline}}{\text{baseline}}$); Lower: The accuracy of accepted paper, described by $\frac{|\widehat{\mathcal{P}}_{@30}\cap{\mathcal{P}}_{@30}|}{|{\mathcal{P}}_{@30}|}$, where ${\mathcal{P}}_{@30}$ is the set of truly top 30% papers (ICLR has acceptance rate  30%).
Q2: The informativeness and robustness tradeoff of the Isotonic Mechanism when there is authors’ perception noise. Our Isotonic Mechanism offers flexibility for practitioners to decide the number of authors from whom they would like to elicit ranking information (i.e., the $L$-strongness of the mechanism). Eliciting from a single author can lead to larger blocks in the partition hence more efficiency, but suffers more risk of miscalibration when the authors have noise perception of the ranking. Our second experiment tests such informativeness vs robustness tradeoff. However, the $L$ parameter can hardly go beyond 1 in the authorship graph of real conference, as the papers that share more one common authors are very rare. This requires us to generate problem instances with carefully structured ownership relations in order for a diverse set of partitions to emerge. We thus use pure simulation data here.
@@
Specifically, in our simulation, the authors perceive score as $R_{i}+\zeta^{j}_{i}$ and control the variance of authors’ perception noise variable $\zeta^{j}_{i}$ in a range from $\{0.1,0.5,1,2\}$ — in a smaller scale than the review noise, as we expect that authors tend to have less noise perception of their papers than the reviewers (roughly conforming to our modeling assumption). Meanwhile, to vary the partition structure, we construct a class of instances whose connections can be characterized by a $7$-level ternary tree with $3^{7}$ items and $(3^{7}-1)/2$ authors. Intuitively, the partition at one extreme is to have only a single block of all items owned by the root node (author), and there is only one author we can elicit. At the other extreme, the partition is to have one block for the set of items owned by each leaf node, and we can elicit from the authors at this leaf node and all its ancestor nodes. Hence, there exist $L$-strong partitions with $L$ from 1 to 7. In Figure 6, we can observe a clear trend of tradeoff: in the case of low noise (blue curve), the large block size is preferred to have more pairwise information, while in the case of high noise (red curve), the smaller (though not necessarily the smallest) block size is preferred to have more authors per block (larger $L$). In cases of medium noise (green or orange curve), the optimal block size is a careful balance of the block size and $L$-strongness. The trend showcases the importance of eliciting from multiple authors, e.g., to enforce the $L$-strong partition constraint, though the parameters should be tailored carefully for the noise level in real-world applications.
Figure 6: The loss (MSE) of Isotonic Mechanisms under different levels of authors’ perception noise with varying number of authors per partition block and different review noise $\sigma=$ 1 (left), 2 (middle), 4 (right).
Q3: Evaluations on the real ICML 2023 data. Our last set of experiments focuses on the ICML dataset which is richer than the ICLR data since this data contains reported ranking from the real-world survey, hence allowing us to conduct almost all real experiments, except that the authors’ report in the survey would not affect the final outcomes of their paper. The performance of Isotonic Mechanisms is presented in Table 5. From these results, we are able to state with high confidence that Isotonic Mechanisms under both partition schemes are capable of reducing the noise in the review process.
Baseline Greedy Random
MSE $1.4267$ $1.3012$ $1.3210$
F-Statistics - $15.828$ $9.911$
P-Value - $7.032\times 10^{-5}$ $1.652\times 10^{-3}$
Table 5: The loss of Isotonic Mechanisms under different partition scheme in the ICML dataset, along with the results of F-test and P-value on whether the loss indeed decreases.
6 Conclusion and Additional Discussions
In this work, we have established the Isotonic Mechanism primarily designed to assist peer review in machine learning conferences, specifically addressing the challenges where multiple authors contribute to various papers. We demonstrate the truthfulness of our mechanism and design a simple polynomial-time algorithm to partition the ownership sets, which achieves a constant ratio approximation of the optimal partition that is computationally hard to find. This facilitates the practical implementation of our mechanism. However, our mechanism is not without limitations, offering avenues for future exploration. We begin with the applicability of the technical assumptions made in our theoretical analysis and point out the potential extensions and relaxations to investigate in future work.
Owner’s Private Information. Assumption 2.1 is a standard private knowledge assumption, commonly adopted in the mechanism design literature (e.g., in auction design it is often assumed that bidders know their values or private signals). Here we assume owners have accurate knowledge of the information we elicit, i.e., relative ranking of their own items. This assumption is not to ask owners to have precise knowledge on ground-truth scores of their items — only agreement on the (much coarser) ranking information suffices. We make a remark on the intrinsic tradeoff when eliciting information from owners. On one hand, eliciting very fine-grained score information can be risky since owners may not have accurate information. On the other hand, eliciting overly coarse information may not be useful any more. As a compromise, we believe the elicitation of ranking information is a reasonable “middle ground” to begin with, though it is an interesting future direction to further explore other possibilities. Additionally, in Section 5 where we evaluate our algorithm on situations with authors’ noisy perception of the true paper ranking, we still observe improvements on score estimation by the Isotonic Mechanisms, indicating that this approach remains statistically effective when owners can supply additional (even somewhat noisy) information not yet included in review scores.
That said, as demonstrated in the following Remark 5, our current mechanism cannot guarantee a truthful Nash equilibrium if some co-owners would “truthfully” report a misleading ranking according to his noisy knowledge of his papers’ relative ranking. The mechanism also ceases to be truthful if an owner can, to some degree, predict the realization of review noise, e.g., due to ex-ante insider information on the reviewers. Therefore, it remains a fundamental design challenge for future work to ensure the authors’ incentive compatibility when they hold outside information or conflicting information with each other. Nevertheless, we believe our current model and its truthfulness result remain valuable, as they encourage coauthors to discuss and refine their perceptions for more accurate rankings. We also conjecture that truthful behavior may be the most robust strategy when authors are uncertain about others’ response model, and we leave a formal analysis to future work.
Remark 5 (On the Strategyproofness of the Isotonic Mechanism).
[…]
$\mathbb{E}_{z}\bigl{[}\,\mathrm{U}(\pi_{2}=(1,3,2))\,\bigr{]}=5.33\quad>\quad 5.25=\mathbb{E}_{z}\bigl{[}\,\mathrm{U}(\pi_{2}=(1,2,3))\,\bigr{]}.$
This example demonstrates that, even in a two-owner, three-item scenario, if one owner’s report is non-truthful (e.g., corrupted by noise), it could be suboptimal for the other owner to report the ranking truthfully.
Noise Structures in Review Scores. The exchangeable noise structure in Assumption 2.2 is also a natural choice and has been widely adopted in mechanism design literature. The typical rationales behind this assumption are as follows: an owner determines the ranking of his paper ex-ante, i.e., before any submission and reviewer assignment have been made, a moment that he does not know how his papers will be reviewed. Thus, the best this owner could assume is the symmetry of review noises for all his papers, which is precisely described by our exchangeable noise assumption. Similar justification on such exchangeable noise assumption is also given in a recent work by [26], though for a completely different mechanism design problem. We point out that our Assumption 2.2 is a strict relaxation of the i.i.d. noise assumption, which is perhaps more often adopted in the peer review literature (See e.g., [1, 25, 38]). Notice that the noise satisfying Assumption 2.2 does not even need to have a zero mean. In addition, if the review score of each paper is averaged over several reviews and the number of reviewers assigned to each paper is treated as an i.i.d. random variable, then the exchangeable noises allow the papers to have different variances.
Finally, an advantage of the exchangeable noise assumption is that it is non-parametric, without the need of modeling the distribution family. If one is willing to assume specific format of the noise distributions, then the exchangeable noise assumption may be further relaxed (e.g., see recent work by Yan et al. [43] using exponential family to model review score distributions). However, without such structural distribution assumption, we conjecture it will be quite difficult (if not impossible) to relax this exchangeable noise assumption in our setup.
Owners’ Utility Structures in Peer Review. The convex utility in Assumption 2.3 may not appear as common as the two assumptions above, hence worth some more discussions here. To begin, it is important to distinguish two traits that affect utilities, the quality and quantity. While utility theory often assumes concavity in the quantity of products due to the law of diminishing return, it is not uncommon to assume utility’s convexity in quality, including the seminal work of Mussa and Rosen [28] on pricing quality-differentiated goods and its numerous follow-up works. [^14] More concretely, for tractability reasons, these papers often normalize utility to be linear (i.e., weakly convex) in quality by re-parameterization, but shift the convexity into cost functions. In our setting, utility convexity in ratings reflects the nature that items receiving higher rating tend to generate significantly more reward. We believe such convex utility is particularly suitable in the context of evaluating research (e.g., conference review), fundamentally due to the “high-risk-high-reward” nature of research, recognized by various funding agencies [42, 7, 12]. A spiritually similar context is in the tournament design [23, 30], where participants are incentivized to exert more effort or take on more risk to compete for the best as the potential rewards are significantly larger at the top; participants’ utilities are commonly assumed to be convex in their ratings (akin to the scores of papers in our setting). In fact, this assumption of convexity in scarcity is naturally consistent with the classic assumption of concavity in quantity. This is because, if a product is hardly accessible (e.g., papers with high scores), or an award is conferred to only a few, we tend to derive more utility — a phenomenon famously known as the “Veblen effects” in behavioral economics and social psychology [40, 41]. In line with the “Veblen effects”, most publication venues have indeed been implementing various metrics to promote (more scarce) high-quality research papers. For instance, with the massive number of accepted papers at AI/ML venues today, these conferences start to distinguish accepted papers with labels such as “poster”, “spotlight” and “oral”; oral papers are promoted with extra resources such as longer presentations and more official media highlights, which naturally lead to much higher utility than an accepted poster paper.
Figure 7: The first two plots based on the ICLR 2022 dataset illustrate the probability of an paper getting accepted as a “poster”, “spotlight”, “oral” w.r.t. its averaged review score. The dashed line denotes the estimated probability from raw data, the smooth line denotes the probability predicted by logistic regression. The last plot illustrates the expected utility curve based on a paper’s chance of receiving different acceptance labels.
[…]
Remark 6 (A case of the “convexifying” owner utilities in peer review).
The strong convexity of utility could result from not only the high-risk-high-reward nature of research but also the artful design of peer review systems. Existing measures include the selection of best papers and selective labels of “oral“ and “spotlight”. We observe that a potential redesign of the review rating metrics can even further convexify owners’ utility. Specifically, the method is to set less level choices for excellent papers (e.g., merging traditional “7:very good” and “8:excellent” to just a single score of “8:excellent”) while use more fine-grained rating for regular papers. Assuming the exchangeable review noises, such kind of designs will make the differences between regular papers wider while the differences between excellent papers narrower. An explicit goal of this mechanism is to better distinguish papers of different qualities at the acceptance borderline. Meanwhile, despite having narrower score differences, the papers of high ratings typically require much less deliberation (e.g., most of them are accepted with probability $>95\%$ at ICLR 2022). More importantly, this mechanism can implicitly encourage the convexity of acceptance curve by, intuitively, “stretching” the middle borderline region of the curve whereas “condensing” its high-score region. Such redesign has already taken place at ICLR 2024, where reviewers can only choose ratings from a list of non-linear scores (“1”,“3”,“5”,“6”,“8”,“10”) — we look forward to its impact on the ongoing review process.
To relax the current utility assumption, one direction is to study the utility functions that are monotone in the adjusted scores with the design of sequential review, despite the downside of prolonged peer review procedures [45]. Meanwhile, an author’s value of a paper could depend on factors such as authorship order, the length of the author’s list, and his relative contributions to the paper [10]. This modeling perspective could more accurately reflect the reality that authors may assign different values to their co-authored submissions. Recognizing the potential heterogeneity of the utility function poses a significant, yet crucial challenge for future work. Another direction is to understand the potential group manipulations or collusion under our mechanism. This is generally a challenging problem in mechanism design, as it requires incentive analysis beyond the unilateral deviation from Nash equilibrium — our Remark 5 shows that truthful behaviors do not form a dominant strategy Nash Equilibrium under the current setup. For future work, one could investigate how defense techniques used in paper assignment [32, Section 4.2] can be adopted to detect and mitigate group manipulations from paper authors.
Lastly, it is important to realize that theory might only take us so far in this field, due to the complex nature of human subjects. We view our work as a theoretical foundation for the initiatives aimed at eliciting self-evaluations and managing incentives in scientific reviews. While empirical results from the OpenRank survey at ICML 2023 are promising [35], further effort is required to implement this method in practice, where various challenges may arise. As an initial step, the Isotonic Mechanism can function as an independent layer atop existing peer review processes, with its calibrated scores serving as reference points to assist expert human decision-making (e.g., helping area chairs to detect anomalies and request additional reviewers). We look forward to more real-world experiments to better understand the potential limitations of the Isotonic Mechanism in practice and drive iterative improvements towards solving the owner-assisted calibration problem.
Acknowledgments
This work was supported in part by NSF through CCF-2303372 and CCF-1934876, ARO through W911NF-23-1-0030, Analytics at Wharton, and Wharton AI and Analytics for Business.
Appendix A Proof of Theorem 1
Before we dive into the proof, we introduce several definitions and lemmas for Theorem 1.
@@ Appendix A Proof of Theorem 1
holds for all convex functions $h:\mathbb{R}\to\mathbb{R}$ if and only if $\bm{a}\succeq\bm{b}$.
Lemma 3.
For any vector $\bm{a},\bm{b}\in\mathbb{R}^{n}$, $\bm{a}^{+}+\bm{b}^{+}\succeq\bm{a}^{+}+\rho\circ\bm{b}^{+}$ for any permutation $\rho$ on $\bm{b}^{+}$, where $\bm{a}^{+}=(a_{(1)},a_{(2)},\dots,a_{(n)})$ and $\bm{b}^{+}=(b_{(1)},b_{(2)},\dots,b_{(n)})$
Proof.
Proof of Lemma 3 It is clear that $(\bm{a}^{+}+\bm{b}^{+})_{(i)}=\bm{a}^{+}_{i}+\bm{b}^{+}_{i}=a_{(i)}+b_{(i)}$. And for any permutation $\rho$, we have $(\rho\circ\bm{b}^{+})_{(i)}=\bm{b}^{+}_{(i)}=b_{(i)}$. Next, we prove the lemma by definition. The equality constraint holds since $\rho$ is a permutation. For any $k<n$, we have
$\sum_{i=1}^{k}(a_{(i)}+b_{(i)})=\sum_{i=1}^{k}a_{(i)}+\sum_{i=1}^{k}(\rho\circ\bm{b}^{+})_{(i)}\geq\sum_{i=1}^{k}(\bm{a}^{+}+\rho\circ\bm{b}^{+})_{(i)}.$
We have proved that $\bm{a}^{+}+\bm{b}^{+}\succeq\bm{a}^{+}+\rho\circ\bm{b}^{+}$ by definition. ∎
Lemma 4.
Proof.
Proof of Lemma 4 We first verify that equality holds. Since $\sum_{i=1}^{n}a_{i}=\sum_{i=1}^{n}a_{i}^{\prime}$,
$\sum_{i=1}^{n}(a_{i}+b_{i})=\sum_{i=1}^{n}(a^{\prime}_{i}+b_{i}).$
$\sum_{i=1}^{k}(a_{i}+b_{i})\geq\sum_{i=1}^{k}(a^{\prime}_{i}+b_{i}).$
Theorem A.1 (Restatement of [36, Theorem 1]).
Proof.
Proof of Theorem 1 We first show that truthful reporting forms a Bayes-Nash equilibrium. Pick any owner and let its utility function be $\mathrm{U}(\widehat{\bm{R}})=\sum_{i=1}^{n}{U}(\widehat{R}_{i})$. Suppose that the other $m-1$ owners are already truth-telling and without loss of generality we can index them from $1$ to $m-1$. With slight abuse of notation, we let $\widehat{\bm{R}}^{j}(\pi)$ be the random variable for the score estimates using owner $j$’s reported ranking $\pi$ w.r.t. the randomness of review noise $\bm{z}$, i.e, $\mathop{\mathbf{Pr}}[\widehat{\bm{R}}^{j}(\pi)=\widehat{\mathrm{R}}^{j}(\pi;\bm{R}+\bm{z})]=\mathop{\mathbf{Pr}}[\bm{z}]$. Suppose that the first $m-1$ estimates use the true ranking and the true ranking $\pi^{\star}$ is the identity permutation for simplicity. It is clear that the item score estimates with all $m-1$ owners’ truthfully reported rankings are all identical,
@@
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\alpha^{m}\widehat{\bm{R}}^{m}(\pi)+(1-\alpha^{m})\widehat{\bm{R}}(\pi^{\star}))\right]=\mathop{\mathbf{E}}_{\bm{z}}\left[\frac{1}{n!}\sum_{\rho}\mathrm{U}\left(\alpha^{m}\widehat{\mathrm{R}}^{m}(\pi;\rho\circ\bm{z})+(1-\alpha^{m})\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})\right)\right].$
@@
Let $\bm{a}^{+}$ be the projection of $\bm{a}$ onto the standard isotonic cone $\{\bm{a}|a_{1}\geq a_{2}\geq\dots\geq a_{n}\}$. Recall that we assume without loss of generality that $R_{1}\geq R_{2}\dots\geq R_{n}$. Following from the coupling argument in the proof of Theorem 1 in [36], we have $\widehat{\mathrm{R}}(\pi^{\star};\bm{z})=(\bm{R}+\bm{z})^{+}$ and $\widehat{\mathrm{R}}(\pi;\bm{z})=\pi^{-1}\circ(\pi\circ\bm{R}+\pi\circ\bm{z})^{+},\forall\bm{z}\in\mathbb{R}^{n}$. So we can rewrite the inequality (4) as,
@@
The first inequality is due to Lemmas 2 and 3, as $\pi^{-1}$ is some permutation on $\bm{b}^{+}=(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}$. The second inequality follows from Lemmas 2, 4 as well as Lemma 2.4 of [36], which shows that $(\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\succeq(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+},\forall\rho$. The third inequality follows from Lemma 2 and 3 where for each distinct permutation $\rho_{j}$, $(1-\alpha^{m})(\bm{R}+\rho_{j}\circ\bm{z})^{+}_{i},\alpha^{m}(\bm{R}+\rho_{j}\circ\bm{z})^{+}_{i}$ are the $j$-th entry of $\bm{a}^{+},\bm{b}^{+}\in\mathbb{R}^{n!}$ and $\pi$ is some permutation on $\bm{b}^{+}$. The equalities are by the definition of the utility function and its linearity.
@@
It now remains to prove that truthful NE gives every owner the highest equilibrium utility among all possible equilibria of the game. Again, we pick arbitrary owner $i$. Under truthful NE, the owner’s utility is $\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{\star}))\right]$, where $\widehat{\bm{R}}(\pi^{\star})$ is the item score estimate under the truthful ranking. We compare it to any (possibly mixed) Nash equilibrium (see Definition 1). Consider any pure strategy profile in this mixed Nash equilibrium, $\bm{\pi}=(\pi^{1},\pi^{2},\dots,\pi^{m})\sim\bm{x}=(x^{1},x^{2},\dots,x^{m})$. This sampled profile of reported ranking corresponds to the weighted average of score estimates, $\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}(\pi^{j})$. We can see that the owner’s expected utility under this pure strategy profile is no larger than that under the truthful strategy profile,
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}\left(\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}(\pi^{j})\right)\right]$ $\leq\sum_{j=1}^{m}\alpha^{j}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{j}))\right]$ (5)
$\leq\mathop{max}_{j\in[m]}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{j}))\right]$
$\leq\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{\star}))\right],$
@@
where the first inequality is due to Jensen’s inequality and linearity of expectation, the second inequality is by the property of arithmetic mean, last inequality is due to Theorem A.1. Hence, the owner’s expected utility under this Nash equilibrium, taking an expectation over the distribution of pure strategy profile, $\mathop{\mathbf{E}}_{\bm{\pi}\sim\bm{x}}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}^{j})\right]\leq\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{\star}))\right]$, is no larger than the expected utility under the truthful NE. ∎
Remark 7.